OpenStudy (anonymous):

Use the distributive property to simplify the expression below: The quantity of 2 times x minus 8 all divided by 2 x − 8 x − 4 2x − 4 x + 2

OpenStudy (debbieg):

Remember that division is the same as muliplication by the reciprocal: \[\Large \dfrac{ 2x-8 }{ 2 }=\dfrac{ 1 }{ 2 }\left(2x-8\right)\] So just distribute the (1/2) through the parens per the distributive property: \(\Large a(b-c)=ab-ac\)